# Where Do We Stand with Implicit Neural Representations? A Technical and Performance Survey

### v1

@@ 3 Experimental Results
In this section, we present a comprehensive comparison of the state-of-the-art implicit neural representation (INR) methods that has publicly available implementations. These methods were evaluated across a diverse set of tasks, ranging from 1D audio reconstruction to various image reconstruction tasks, including denoising, super-resolution, and CT reconstruction, as well as a 3D column occupancy task. This broad evaluation aims to highlight the strengths and limitations of each method in different application domains, providing a holistic view of their capabilities and performance across multiple dimensions.
For a fair comparison of all implemented INR methods, we implemented each approach using the default hyperparameters recommended in the respective papers. All models were trained and tested on an RTX 4070 GPU with 8 GB of RAM.
@@ 3.2.1 CT Reconstruction
$\mathcal{L}=\left\|\text{sinogram}(output)-\text{sinogram}(truth)\right\|$ (20)
@@
Both Figure 4 and Table 1 indicate that the reconstruction quality improves for all methods as the number of projections increases, which aligns with the general expectation that more projections provide richer information, leading to more accurate and detailed reconstructions. Observing from Table 1, Incode demonstrates superior performance, achieving the highest PSNR (Peak Signal-to-Noise Ratio) and SSIM (Structural Similarity Index Measure) values in all scenarios, particularly excelling at higher projection counts with scores of 34.76 dB for PSNR and 0.953 for SSIM at 300 projections. At lower projection counts, Fr stands out, delivering the best results with a PSNR of 29.14 dB and SSIM of 0.852 at 50 projections, and further improving to 31.18 (dB) PSNR and 0.917 SSIM at 100 projections, making it highly effective with limited data. Finer performs competitively as the number of projections increases, while Wire and Mfn consistently lag behind, exhibiting the lowest scores across all scenarios. Moderate performers like Fourier Features and Gauss show adequate results in low to medium projection settings but are outperformed by Incode and Fr at higher counts. These findings underscore the strengths and limitations of each method, highlighting Incode as the most robust approach for high-quality CT reconstruction, especially when more projection data is available.
@@ 3.2.2 Image Denoising
We report the results in Figure 5 and Table 2.
Siren also performs well, with a reasonable trade-off between accuracy (PSNR 28.60 dB) and computational speed (482 seconds). On the other hand, Fourier Features, while the fastest (403 seconds), sacrifices reconstruction quality, evident in its lower PSNR (27.16 dB) and noisier output. Methods like Wire, Gauss, and Mfn fall behind in both visual quality and PSNR, with their outputs appearing either over-smoothed or retaining significant noise, making them less competitive for high-quality denoising tasks.
@@ 3.2.3 Single Image Super Resolution
While observing from the visualisation results, in Figure 678, and 9, the conclusion coincide with the table result with more detailed features. From the zoom in view of Figure 67 for the lower resolution task, we observe Incode and Finer successfully reconstructed the detailed fur at the animal’s ear and the mustache. However, as the scaling increases, Incode starts to show color artifacts in the reconstruction result, while Finer oversmooths the details. This oversmoothing is likely a result of Finer’s lack of spatial compactness, which prevents it from preserving fine features at higher resolutions. In fact several other methods has shown artifacts with different features when the super resolution scale increases. For example, the Fourier based methods that lack of compact frequency and spatial representations shows undesirable repetitive patterns at finer scales, which Fr exhibits twisting pattern blobs, and Fourier Features shows the grid like artifacts. Similarly, noise artifacts become more alleviated in the reconstruction results of Wire, Mfn, Gauss, and Siren all of which lack the task-specific parameters necessary for tuning to different levels of super-resolution scales. This adaptability proves to be beneficial, as both the visualisations and the metrics show that Incode, Finer, and Fr which has this task-specific parameters perform better across different scales.
@@
Table 4: 3D Occupancy Metrics
@@ 4 Discussion & Future Work
Finally, focusing on methods that efficiently integrate both spatial and frequency compactness while reducing computational demands should be a priority for future research. Such advancements would make INR models more accessible for real-time applications while maintaining high fidelity in complex tasks like image and audio synthesis.
Acknowledgment
YC is funded by an AstraZeneca studentship and a Google studentship. CWC is funded by Swiss National Science Foundation (SNSF). CBS acknowledges support from the Philip Leverhulme Prize, the Royal Society Wolfson Fellowship, the EPSRC advanced career fellowship EP/V029428/1, EPSRC grants EP/S026045/1 and EP/T003553/1, EP/N014588/1, EP/T017961/1, the Wellcome Innovator Awards 215733/Z/19/Z and 221633/Z/20/Z, the European Union Horizon 2020 research and innovation programme under the Marie Skodowska-Curie grant agreement No. 777826 NoMADS, the Cantab Capital Institute for the Mathematics of Information and the Alan Turing Institute. AIAR acknowledges support from CMIH (EP/T017961/1) and CCIMI, University of Cambridge.

### v2

@@ 3 Experimental Results
In this section, we present a comprehensive comparison of the state-of-the-art implicit neural representation (INR) methods that have publicly available implementations. These methods were evaluated across a diverse set of tasks, ranging from 1D audio reconstruction to various image reconstruction tasks, including denoising, super-resolution, and CT reconstruction, as well as a 3D column occupancy task. This broad evaluation aims to highlight the strengths and limitations of each method in different application domains, providing a holistic view of their capabilities and performance across multiple dimensions.
For a fair comparison of all implemented INR methods, we implemented each approach using the default hyperparameters, per tasks, recommended in the respective papers. All models were trained and tested on an RTX 4070 GPU with 8 GB of RAM.
@@ 3.2.1 CT Reconstruction
$\mathcal{L}=\left\|\text{sinogram}(output)-\text{sinogram}(truth)\right\|,$ (20)
@@
Both Figure 5 and Table 2 indicate that the reconstruction quality improves for all methods as the number of projections increases, which aligns with the general expectation that more projections provide richer information, leading to more accurate and detailed reconstructions. Observing from Table 2, Incode demonstrates superior performance, achieving the highest PSNR (Peak Signal-to-Noise Ratio) and SSIM (Structural Similarity Index Measure) values in all scenarios, particularly excelling at higher projection counts with scores of 34.76 dB for PSNR and 0.953 for SSIM at 300 projections. At lower projection counts, Fr stands out, delivering the best results with a PSNR of 29.14 and SSIM of 0.852 at 50 projections, and further improving to 31.18 PSNR and 0.917 SSIM at 100 projections, making it highly effective with limited data. Finer performs competitively as the number of projections increases, while Wire and Mfn consistently lag behind, exhibiting the lowest scores across all scenarios. Moderate performers like Fourier Features and Gauss show adequate results in low to medium projection settings but are outperformed by Incode and Fr at higher counts. These findings underscore the strengths and limitations of each method, highlighting Incode as the most robust approach for high-quality CT reconstruction, especially when more projection data is available.
@@ 3.2.2 Image Denoising
Siren also performs well, with a reasonable trade-off between accuracy (PSNR 28.60) and computational speed (482 seconds). On the other hand, Fourier Features, while the fastest (403 seconds), sacrifices reconstruction quality, evident in its lower PSNR (27.16) and noisier output. Methods like Wire, Gauss, and Mfn fall behind in both visual quality and PSNR, with their outputs appearing either over-smoothed or retaining significant noise, making them less competitive for high-quality denoising tasks.
@@ 3.2.3 Single Image Super Resolution
While observing from the visualisation results, in Figure 789, and 10, the conclusion coincide with the table result with more detailed features. From the zoom in view of Figure 78 for the lower resolution task, we observe Incode and Finer successfully reconstructed the detailed fur at the animal’s ear and the mustache. However, as the scaling increases, Incode starts to show color artifacts in the reconstruction result, while Finer oversmooths the details. This oversmoothing is likely a result of Finer’s lack of spatial compactness, which prevents it from preserving fine features at higher resolutions. In fact several other methods has shown artifacts with different features when the super resolution scale increases. For example, the Fourier based methods that lack of compact frequency and spatial representations shows undesirable repetitive patterns at finer scales, which Fr exhibits twisting pattern blobs, and Fourier Features shows the grid like artifacts. Similarly, noise artifacts become more alleviated in the reconstruction results of Wire, Mfn, Gauss, and Siren all of which lack the task-specific parameters necessary for tuning to different levels of super-resolution scales. This adaptability proves to be beneficial, as both the visualisations and the metrics show that Incode, Finer, and Fr which has this task-specific parameters perform better across different scales.
@@
Table 5: The metric comparison of 3D occupancy task across the 8 INR methods.
@@ 4 Discussion & Future Work
Finally, focusing on methods that efficiently integrate both spatial and frequency compactness while reducing computational demands should be a priority for future research. Such advancements would make INR models more accessible for real-time applications while maintaining high fidelity in complex tasks like image and audio synthesis. While the surveyed methods provide a comprehensive and well-structured taxonomy of implicit neural representations (INRs), there are several areas present opportunities for deeper exploration to enhance our understanding of this domain. For instance, while the interplay between activation functions and position encoding is acknowledged, further investigation into how these components synergise could reveal additional avenues for optimisation. Similarly, potential approaches beyond the proposed taxonomy, such as hybrid models that integrate INRs with explicit representations, offer a promising direction for enhancing performance in tasks requiring both adaptability and precision, such as 3D reconstruction. Furthermore, the influence of hyperparameters, including encoding frequencies, network architectures, and activation scaling, remains a pivotal factor in fine-tuning method effectiveness across various data modalities. By building on the insights provided in this survey, future work can explore these areas to refine and expand the utility of INR techniques in a broader range of applications. The development of implicit neural representations (INRs) has the potential to transform various domains positively. However, it is crucial to consider ethical implications and ensure their responsible use to avoid unintended societal consequences.